OpenStudy (anonymous):

How do i graph this function? y = -(x+4) - 1 also show how to do it

OpenStudy (shamil98):

First put the equation in slope-intercept form. y = -x -4 -1 y =-x -5 Slope is -1/1 and the y intercept is -5.

OpenStudy (nubeer):

lol too much work :P

OpenStudy (debbieg):

Surely, if she is asking how to do it here, she is expected to be able to do it without technology, don't you think?

OpenStudy (shamil98):

Plot the y intercept on the graph and make a line using the given slope, (-1/1). down 1 right 1 or up 1 left 1.

OpenStudy (debbieg):

If you are doing transformations of functions, then it might be expected that you would do it with transformations. This is a transformation of the equation y=x. But if not, then @shamil98 's method is perfect.

OpenStudy (anonymous):

were supposed to make a parabola lol im so confused but would i plot -1 and then do -1/1

OpenStudy (shamil98):

There are no mentions of trigonometric functions (e.g. sin , cos, tan, cotan, etc.) so I assumed it would be this method?.. but if it was it would me different with phase shifts and vertical shifts.

OpenStudy (debbieg):

Then you gave us the wrong equation, lol. Is the stuff in the ( )'s squared??

OpenStudy (anonymous):

yeah its squared thats exactly how the problem is

OpenStudy (shamil98):

Oh, a parabola. My bad.

OpenStudy (debbieg):

@shamil98 I teach transformations in algebra 2, no trig required.

OpenStudy (debbieg):

If that's the equation you are given, that is NOT a parabola. If you are given \(y = -(x+4)^2 - 1\) then THAT is a parabola.

OpenStudy (shamil98):

The squared part should be shown as ^2 , for future reference , helps us helpers know what to do. lol.

OpenStudy (anonymous):

yeah sorry its the one you just said @DebbieG

OpenStudy (debbieg):

ok, if it's \(y = -(x+4)^2 - 1\), then again, it kind of depends on what you've done in your class. If you have done transformations of functions, I would explain it one way. If you haven't but are just being introduced to graphing quadratic equations, then I would explain it a bit differently... so I'll assume the latter.

OpenStudy (debbieg):

\(y = -(x+4)^2 - 1\) is called "vertex form", because you can read the vertex right off the equation: \(y = a(x-h)^2 +k\) has vertex (h,k) Does that look familiar? So what is the vertex of YOUR equation, \(y = -(x+4)^2 - 1\) ?

OpenStudy (anonymous):

xy i think ??

OpenStudy (debbieg):

Once you have the vertex, you can look at the sign of a to decide it it opens up or opens down. If a>0, opens up if a<0, opens down.

OpenStudy (debbieg):

The vertex is a POINT, so you should have a coordinate pair, (x, y) The question is, what is x, and what is y?

OpenStudy (debbieg):

Look at \(y = a(x-h)^2 +k\) and compare it to your equation, \(y = -(x+4)^2 - 1\) Find h and k. this might help: \(y = -(x-\color{red}{(-4)})^2 +\color{red}{(- 1)}\) \(y = a(x-\color{red}h)^2 +\color{red}k\)

OpenStudy (anonymous):

so would you fill in the x and y slots with the -4 and -1

OpenStudy (debbieg):

Yes, that gives you the point that is the vertex. (-4, -1) is the vertex. Does this sound familiar? This should be what you are covering in this class, given that you are asked this problem

OpenStudy (debbieg):

You know what the "vertex" of a parabola is, right?

OpenStudy (anonymous):

well i missed the last 3 days so im behind and the teacher didnt really help me catch up so i dont know most of this lol

OpenStudy (debbieg):

OK, well, at a minimum you should sit down and read your textbook before trying to jump into the problems; and get class notes from someone. But we can move forward, as long as you are understanding, so far.

OpenStudy (debbieg):

OK, so we have the vertex. Now, which direction does this parabola open - up or down? (I talked about that above, scroll up if you need to read it again)

OpenStudy (anonymous):

down

OpenStudy (debbieg):

Very good. so we have: \(y = -(x+4)^2 - 1\) has vertex at (-4,-1) and opens downward. Now we want to see if we can figure out a couple other things to help us with the graph. How about the y-intercept? How do we find that?

OpenStudy (debbieg):

Just like for a line or another graph, the y-intercept happens when ___ =0? What goes in the ___ ? :)

OpenStudy (anonymous):

when x = 0 i think right ? im sorry im terrible at math

OpenStudy (debbieg):

Yes, when x=0. Don't ever say that you're terrible at math... you've said that your whole life and now you're convinced of it, aren't you? :) Now we want to find the value of that y-intercept, and we know it happens when x=0. So we just plug in x=0 into our equation: \(y = -(x+4)^2 - 1\) so we have \(y = -(0+4)^2 - 1=???\)

OpenStudy (debbieg):

can you evaluate that and tell me what that y-value will be?

OpenStudy (anonymous):

15 ?

OpenStudy (debbieg):

oops,close, but the - sign in front of the ( )'s happens AFTER you square the 4, right?

OpenStudy (debbieg):

\(y = -(4)^2 - 1= -16 - 1=??\)

OpenStudy (anonymous):

oh so -17

OpenStudy (debbieg):

See, it ISN'T \((-4)^2\), it's \(-(4)^2\), that's the difference. RIGHT, -17.

OpenStudy (debbieg):

OK, now we have the vertex, we know it opens downward and we know that the y-intercept is at -17. That's quite a bit of info so far: 

OpenStudy (debbieg):

Now, the next thing that we would usually do is to find the x-intercepts... which is where y=0 (where the graph crosses the x-axis. Only, with this function, there is one problem with that. Do you see what it is? Why am I not going to look for x-intercepts?

OpenStudy (anonymous):

i have no clue lol

OpenStudy (debbieg):

lol.. ok, look at some different parabolas:  Notice that one of the downward opening ones has x-intercepts, the other doesn't. Notice that one of the upward opening ones has x-intercepts, the other doesn't. Do you see what makes the difference?